The ratio of a jacket to the original price is 2:3. The original price is $65. What is the sale price?
Answer by Alan3354(30993)

A ratio should be between similar items, such as a price to another price.
Not a jacket to a price.
------------------------
If you mean the sale price to the OP is 2:3 then,
SP:OP::2:3
3SP = 2OP = 2*65 = 130
SP = 130/3
SP = $43.33